Summary

Overall a productive night. Seeing 0.7"-1.3". Observed one OB in UMaII, one OB in Herc, six OBs in Bootes and three OBs in EGS. Lost the last hour of the night due to power failure and slow recovery from it.

Details

Observers: Xiaohui Fan, Ben Weiner, Linhua Jiang

Supporting Astronomer: Marco Pedani

Operator: David Gonzalez Huerta

All Times UT

0115 reboot CMU

0125 25Bias_Bino

0202 SkyFlat_UY_5

0208 SkyFlat_BR_5; red: only one chip read out; blue is fine. Tunned out for some unknown reason, only chip2 was set to read out in the SkyFlat_BR_5.ob. It affected last night's R sky flat as well.

0220 SkyFlat_Vr_5

0250 going to UMaII, chasing focus

0330 UMaII_offset12_Vr; first image elongated in blue, likely bad guiding? red OK, 1.3" in red. Second image OK. both 1.3". improving to about 1".

0425 NGC_3368_10step; 1" in both V and R; John did some on-the-fly active optics to improve image quality.

0515 Bootes_48; a few images lose guiding in blue.

0545 Bootes_59; seeing still about 1"

0620 Bootes_61

0655 Bootes_66; seeing 0.8-0.9"

0730 Bootes_68; seeing 0.8-0.9"

0808 EGS_Y_3a; seeing 0.7-0.8 in red

0843 EGS_Y_3b; seeing 0.7-0.8 in red

0917 EGS_Y_4a; seeing 0.9-1.1 in red

1000 Bootes_70; first Y image elongated, others good; seeing 0.7-0.8 in red

1049 Herc_1; power down, stop exposure; problem with recovering from power failure; we lost more than 1 hours

1228 sky flats SkyFlat_BR_5, SkyFlat_Vr_5, and SkyFlat_UY_5; it's a little late, so some images are saturated

-- XiaohuiFan - 08 Apr 2008
Topic revision: r5 - 14 Apr 2008, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback